40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 13 15:31:09 2012, steve, Update, General, larger optical tables at the ends ETMYtable.jpgETMY4X3.jpg
    Reply  Mon Aug 13 18:00:30 2012, jamie, Update, General, larger optical tables at the ends 
       Reply  Tue Aug 14 15:58:44 2012, Jenne, Update, General, Translation to English: larger optical tables at the ends 
    Reply  Wed Aug 15 16:01:47 2012, steve, Update, General, larger optical tables at the ends ? bottleneck.jpg
       Reply  Wed Aug 22 11:41:20 2012, steve, Update, General, grout plate for optical table at the ends IMG_1561.JPG08231201.PDF
          Reply  Thu Sep 27 19:03:06 2012, Manasa, Update, 40m Upgrading, New ETMX layout - Version 1.0 ETMX_OldNew.png
             Reply  Thu Sep 27 20:05:57 2012, Koji, Update, 40m Upgrading, New ETMX layout - Version 1.0 
                Reply  Fri Sep 28 07:49:30 2012, Steve, Update, 40m Upgrading, New ETMX layout - Version 1.0 
Message ID: 7450     Entry time: Thu Sep 27 20:05:57 2012     In reply to: 7449     Reply to this: 7451
Author: Koji 
Type: Update 
Category: 40m Upgrading 
Subject: New ETMX layout - Version 1.0 

1. I wonder how the mode profiling/matching was considered in the new layout.

I can see the distances between the components and lenses are largely different from the old ones.

This is OK if you plan to go through a new mode matching solution with new lenses.
But it takes
 a certain amount of time.

Note that we don't care the distance after the last lens as the Rayleigh range there is supposed
to be long enough to allow this kind of change.

2. The huge frustration of the green alignment in the old setup was caused by the 3D beam steering
at the last two 2" mirrors. i.e. the beam elevation on the table does not match with the beam elevation of the cavity.
In order to avoid this, I suggest you to use three 45 deg 2" mirrors instead of two. In fact these mirrors are supposed to be used at 45deg incidence!

3. The incident green beam and the transmitted IR beam should share a same path as they
 share a same cavity mode.
This means that you should use a harmonic separator for the transmitted light pick-off.

4. Use the harmonic separator for the fiber path too. Get the mirror spec from Jamie.

5. Since the optical window on the chamber has a wedge angle, the beam paths are not straightforward.
The cavity beams can't be moved as they are constrained by the arm cavity.
Probably there is almost no freedom to move even for the oplev beams.
It would be safe just to follow the old positions and angles on the window.
Make sure the beam on the drawing is realistic. The angles of the oplev beams in the old setup look strange.

Is there a possibility to replace the optical window so that it has an AR for 532 and 1064 at least???

6. I wonder if the rejected beam by the Farady have a realistic angle or not. Check it with the old setup.
It is definitely better to have a steering mirrror and a lens before the refl PD.

7. The IR QPD and trans PD are intended to be used for the low and high power detection.
I forgot which is which. So check the range of them and think about the power distribution.

8. We should have separated CCDs for IR Trans and Green Refl.
We had a terrible ghost green beam on the IR trans CCD.
Thus, think about the amount of ghost reflection and consider filtering if necessary.

ELOG V3.1.3-